OpenStudy (anonymous):

Alright, so thanks to a couple of wonderful people - I DO have somewhat of a better understanding of systems of equations with 3 variables. BUT these ones threw me off. Can someone please help? (Elimination/Substitution) x-z=5 y+3z=12 2x=y=7 AND x+y=5 3x+z=2 4y-z=8

OpenStudy (mathstudent55):

What's the third equation of the first system?

OpenStudy (mathstudent55):

Let's start with the second system and use substitution. Solve the first equation for y: x + y = 5 y = 5 - x Now solve the second equation for z: 3x + z = 2 z = 2 - 3x Now that you know what y and z are interms of x, plug those results into the third equation: 4y - z = 8 4(5 - x) - (2 - 3x) = 8 Solve for x: 20 - 4x - 2 + 3x = 8 18 -x = 8 -x = -10 x = 10 Now substitute x into the original first equation: x + y = 5 10 + y = 5 y = -5 Now substitue x into the original second equation: 3x + z = 2 3(10) + z = 2 30 + z = 2 z = -28 The solution is: x = 10, y = -5, z = -28

OpenStudy (anonymous):

oh sorry! The first set, 3rd equation is 2x+y=7

OpenStudy (mathstudent55):

Since the "+" character and the "=" character are on the same key on the keyboard, I'm going to assume that the third equation of the first system is: 2x + y = 7. We have this system then: x - z = 5 y + 3z = 12 2x + y = 7

OpenStudy (mathstudent55):

Yes, I figured it out.

OpenStudy (anonymous):

Thank you so much! You have no idea how helpful you are!

OpenStudy (mathstudent55):

Let's use elimination on the first system. Multiply the first equation by -2 and add to the third equation: -2x + 2z = -10 2x + y = 7 ------------------- y + 2z = -3 Notice now that the original second equation also has only the two variables y and z. Rewrite this new equation and below it, multiply the original second equation by -1. Then add them: y + 2z = -3 -y - 3z = -12 ______________ -z = -15 z = 15 Now substitute value of z into first original equation: x - z = 5 x - 15 = 5 x = 20 Now substitute z = 15 into second original equation: y + 3z = 12 y + 3(15) = 12 y + 45 = 12 y = -33 Solution is: x = 20, y = -33, z = 15

OpenStudy (anonymous):

You are wonderful! Seriously thank you!